OpenStudy (anonymous):

Hanna makes her own furniture polish using 30% vinegar and 70% olive oil. Write and solve an equation to determine how much total solution she made if she used fourteen tablespoons of olive oil. Show your work.

OpenStudy (anonymous):

@iambatman @King.Void. @mathstudent55 @Gabylovesyou @Abhisar @Luigi0210 @DarkMoonZ @RAINIAR

OpenStudy (anonymous):

PLease help i only have 3 minutes!!!!

OpenStudy (anonymous):

As soon as I leave it will automatically submit itself! I have to eat then leave please!

OpenStudy (anonymous):

70% olive oil is used and the quantity of olive oil is 14 tablespoons therefore, 70% of the solution is 14 tablespoons Let the total solution be x (70/100)x = 14 x = 20

OpenStudy (anonymous):

But there are 14 tablespoons not 9

OpenStudy (anonymous):

i deleted my first response because i read the question wrong. I corrected myself in a new response, just refresh the page, it should be there.

OpenStudy (anonymous):

@osamabinswaggin

OpenStudy (anonymous):

There was a problem I did exaclty like this one and got that answer with nine tablespoons but this is 14 tablespoons

OpenStudy (anonymous):

The problem said that 70% of the solution was olive oil.14 spoons of olive oil was used. This means that the remainder is vinegar. 30% of the solution is vinegar with is "x" amount of spoons.

OpenStudy (anonymous):

Let the total solution be x (70/100)x = 14

OpenStudy (anonymous):

It could be any answer except for 20 because I got twenty on the last one and the last one was different

OpenStudy (anonymous):

20 is the total amount of solution

OpenStudy (anonymous):

20-14 = amount of vinegar used

OpenStudy (anonymous):

so 6 is the answer

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

6 is 30% of 20 14 is 70% of 20

OpenStudy (anonymous):

well that sucks because I already submitted it :(

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You're welcome, feel free to tag me in anything else you need help with